A car and a bus set out at 2 pm from the same point, headed in the same direction. The average speed of the car is 30 mph less than twice the average speed of the bus. In 2 hrs, the car is 30 miles ahead of the bus. Find the rate of the car.
Answer by ikleyn(52798)

.
A car and a bus set out at 2 pm from the same point, headed in the same direction.
The average speed of the car is 30 mph less than twice the average speed of the bus.
In 2 hrs, the car is 30 miles ahead of the bus. Find the rate of the car.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

The first equation is

C = 2B - 30,     (1)   where C is the car speed, B is the bus speed.

The second equation is 

2C - 2B = 30.    (2)


Rewrite the equations in this form

 C - 2B = -30,    (1')
2C - 2B =  30.    (2')

Distract the equation (1') from equation (2') (both sides). You will get

2C - C = 30 - (-30),

C = 60.

So, you just found the speed of the car it is 60 mph.

From this point please complete the solution on your own.